The perimeter of a triangle is 51 cm. The longest side is twice the length of the shortest side. The third side is 3 cm longer than the shortest side. How long is each side? How do i write an equation that represents the problem and then solve it?
Answer by josgarithmetic(39617)

Assign variables to everything and then use the description to refine what those variables are.
p = perimeter and p=51.
x, y, z are the sides of the triangle.
x = shortest
y = medium
z = longest
-
z=2x.
y=x+3.
Notice how y and z both are in terms of x. What then have you for x, y, z?
-
x = shortest side
x+3 = medium side
2x = longest side

Meaning of perimeter means you have (x)+(x+3)+(2x)=p